] → [a, 시 be continuous. Prove that there exists c E [a,b (3) Let f : [a, such that f(c) = c i.e f has a fixed point
Answer #1

Consider g(x)=f(x)−x.

f(a)≥a so g(a)=f(a)−a≥0.

f(b)≤b so g(b)=f(b)−b≤0.

By the Intermediate Value Theorem, since g is continuous and 0∈[g(b),g(a)] there exists c∈[a,b] such that g(c)=f(c)−c=0, so f(c)=c for some c∈[a,b].